# BoostStep: Boosting mathematical capability of Large Language Models via improved single-step reasoning

### v1

@@ 3.4 Step-Level Example Guidance in MCTS
Detailed ablation experiments in Sec. 4.5 demonstrate that both strategies contribute positively to the overall reasoning quality of Monte Carlo Tree Search (MCTS) methods.
@@ 4.1 Experiment setting
Evaluation Benchmark.  We tested our approach on several challenging open-source mathematical benchmarks, including MATH500 Hendrycks et al. (2021), AQuA Ling et al. (2017), OlympiadBench-TO He et al. (2024) and MATHBench Liu et al. (2024a) College-level and High-level tasks. In addition, we manually collected a selection of problems from the AMC-10 and AMC-12 competitions to serve as even more challenging benchmarks. This is available at https://github.com/beichenzbc/BoostStep . To simulate benchmarks with lower similarity to the example problem bank, we also conducted tests on MathVision Wang et al. (2024) and MathVerse Zhang et al. (2025), highly challenging multi-modal math benchmarks
@@
Hyper-Parameters. The temperature value is 0 in all the experiments except for Monte Carlo Tree Search, which needs some random sampling to generate different reasoning candidates and the temperature value for MCTS is set 0.3. The reference rejection threshold is 0.7.
Prompt. Apart from some necessary guidance like step-level reasoning, we ensured that the prompts for each method were as similar as possible to make the comparison fairer. The specific prompts is listed in the supplementary materials.
@@ 4.2 Comparing to Problem-Level ICL
Table 1: A comparison of different in-context learning strategies on different benchmarks on GPT-4o and Qwen2.5-Math-72B-Instruct. The example problem bank is constructed from PRM800K, so MATH500 is an in-domain benchmark while others are all out-domain benchmarks. Best results are in bold.
@@
We conduct a rigorous comparison of the proposed step-level in-context learning approach and traditional problem-level few-shot learning across multiple benchmarks and base models. Specifically, for traditional problem-level few-shot learning, we set the shot num to 4, which is a common setting for evaluation. The results are presented in Tab. 1. Our step-level In Context-Learning achieves improvements across various mathematical benchmarks, with particularly notable enhancements observed on the relatively more challenging AMC problems.
@@
few-shot 28.7 53.2
Ours 35.2 54.2
Table 3: Experiments on the sensitivity of the similarity between the question and the example problem bank. R_t indicates that the examples are the t_th similar for different method without any rejection strategy. Given a less similar example, our method suffers an 2.26% performance loss, which is much lower comparing to few-shot learning(4.4%).
@@
Sinece the sensitivity to example problem banks is a critical factor limiting the generalizability of in-context learning, we have designed an additional experiment to assess the sensitivity of both step-level and problem-level in-context learning approaches. Specifically, we also test different in-context learning methods on multi-modal mathematical benchmarks including MathVision Wang et al. (2024) and MathVerse Zhang et al. (2025), which has much lower similarity with our example problem bank. The results are shown in Tab. 2. Problem-level few-shot learning not only fails to enhance reasoning performance but can also negatively impact reasoning quality, while ours continues to achieve appreciable improvements.
Moreover, we also manually decrease the similarity between the examples and the problems by selecting the t_th similar example during reasoning. The result is shown in Tab. 3. We can observe that traditional problem-level in-context learning suffers from a severe decrease and is even worse than 0-shot learning when t is larger than 4. In contrast, our method doesn’t not show a significant decline.
These experiments suggest that traditional problem-level few-shot learning requires a high similarity within the example problem bank, which limits its generalizability. Our approach, in contrast, refines the granularity to the step level. Since different problems can still contain similar steps, our method consistently achieves improvements even the examples are dissimilar, demonstrating better general applicability.
@@ 4.3 Construction of Example Problem Bank
To better align with the steps in reasoning, we propose constructing a step-level question bank based on the reasoning content rather than grammatical divisions. To prove our assumption, we compare our approach with a commonly used strategy that constructs steps based on grammatical segmentation, using periods ’.’ as the delimiter, on the same dataset PRM800K and under identical conditions. Results are presented in tab. 4. Our method largely outperforms those using periods as delimiter.
@@ 4.4 Comparison of Retrieving Strategies
The key factor in determining the effectiveness of in-context learning lies in the relevance of the retrieved examples. At the finer-grained step level, designing an appropriate retrieval strategy becomes even more crucial and challenging. Therefore, we propose the first-try strategy, which involves understanding what the model currently needs to reason about using a first attempt, and then searching the problem set for similar steps to guide the model in fully outputting the current step. To validate the effectiveness of this method, we compare it with several other strategies mentioned in Sec.3.3, retrieving by the entire reasoning path $s_{i-1},s_{i-2},\ldots,s_{1},q$ or only by the immediately preceding step $s_{i-1}$.
@@
Table 5: Comparison on different retrieval strategies in step-level in-context learning. The base model is GPT-4o and all the prompts are the same. ’Path’ represents retrieving by the reasoning path including all previous step $s_{i-1},s_{i-2},\ldots,s_{1}$ and question $q$, while ’Pre-Step’ represents retrieving by only the immediately preceding step $s_{i-1}$. Best results are in bold.
@@
4.5 Example-guided Monte Carlo Tree Search
Table 6: A detailed ablation on incorporating retrieving similar steps to provide fine-grained guidance during the reasoning and verifying phases of Monte Carlo Tree Search (MCTS) methods. Base models are GPT-4o and prompts are the same. Best results are in bold.
@@ 4.5 Example-guided Monte Carlo Tree Search
As discussed above, the reasoning capability of generating model and the verifying capability of critic model are two core factors of Monte Carlo Tree Search methods, and our strategy can enhance the reasoning quality of MCTS in both ways. On one hand, it can improve the accuracy of generating candidate nodes using the previously mentioned first-try strategy when reasoning nodes are generated. On the other hand, it can increase the accuracy of evaluation by introducing similar examples during critic model assessments and therefore ensures that the correct reasoning nodes are more likely to be preserved. These two strategies can be decoupled, allowing us to demonstrate the effectiveness of each component through ablation studies.
In the setup for MCTS, we utilize GPT-4o as reasoning model and GPT-4o-mini as Process-supervised Reward Model. Moreover, we adopted the Pairwise Preference Reward Model (PPRM) configuration Zhang et al. (2024b)which reduces variability and ensures a more robust evaluation. Detailed setting will be listed in appendix.
Tab. 6 presents the results of integrating in-context learning into the reasoning and evaluation phases of the Monte Carlo Tree Search methods. The results indicate that introducing examples enhances the verifying and reasoning capabilities of MCTS methods, therefore both approaches contribute to the improvement of overall reasoning performance through providing similar groundtruth steps. Specifically, introducing example steps in reasoning phases will bring more benefits to the overall reasoning quality.
@@ 4.6 Case Study
Figure 4: A specific example of adjusting reasoning during real-time inference through step-level in-context learning. The first-try uses a wrong equation while the retrieving example step guides the model to use the correct equation and get the correct conclusion.
Here we demonstrate a specific example of how our step-level in-context learning boosts step-level reasoning. Given the question, we first let the model to have a first-try on step one. Unfortunately, because the model is unfamiliar with trigonometric functions, it makes an error on tangent sum formula, therefore leading to a wrong step. However, we can get a rough idea on what the model wants to calculate at this step according to the first try. Then, we find a similar step which correctly leverage tangent sum formula in the step-level example problem bank. Therefore, with the guidance provided, the model correctly applied the tangent sum formula during the second reasoning attempt and arrived at the correct answer.
@@ 5 Conclusion
We propose step-level in-context learning, which provides real-time, fine-grained guidance during the reasoning process by searching for similar steps according to the first-try reasoning attempt. This approach improves the model’s reasoning capabilities and reduces the dependency on the similarity of example problem set, thereby increasing the generalizability of in-context learning. Moreover, our method can also enhance the reasoning and evaluation capability of Monte Carlo Tree Search (MCTS) by introducing similar steps in reasoning and verifying phases respectively, thereby improving the overall reasoning correctness.
Potential Limitations. Currently, our example problem bank is entirely sourced from PRM800k, resulting in a relatively homogeneous distribution of example problems and example steps. Furthermore, the TF-IDF retriever is based on modeling language term frequency directly and lacks an understanding of mathematical content, which limits its retrieval capabilities on math problems.
@@ A.3 Benchmarks
We tested our approach on several challenging open-source mathematical benchmarks, including MATH500 Hendrycks et al. (2021), AQuA Ling et al. (2017), OlympiadBench-TO He et al. (2024) and MATHBench Liu et al. (2024a). Specifically, we use the Olympiad-TO (text-only) subset of OlympiadBench, and the application problems in college-level and high-level difficulty.
@@
Appendix B Detailed Setup for MCTS
In the setup for MCTS, we utilize GPT-4o as the reasoning model and employ GPT-4o-mini as the Process-supervised Reward Model (PRM). For the PRM, we adopted the Pairwise Preference Reward Model (PPRM) configuration Zhang et al. (2024b). Specifically, PPRM transforms the absolute rewards calculation into preference predictions between solutions to calculate rewards. This approach reduces the variability associated with scoring characteristics and thus leads to a more robust and consistent evaluation of different solutions.
The complete reasoning process of MCTS in our experiment is as follows: we start with the target problem as the root node and obtain two initial solution steps through sampling to serve as the two initial parent nodes. In each step-level reasoning phase, we expand these two parent nodes through sampling, generating four candidate child nodes. Using the PPRM, we select the two child nodes with higher confidence to become the parent nodes for the next step of reasoning. This process continues until both candidate nodes have completed their reasoning paths, resulting in the final answers. Finally, PPRM is used to select the ultimate answer from these two reasoning paths.

### v2

@@ 3.4 Step-Level Guidance in Tree Search
Detailed ablation experiments in Sec. 4.5 demonstrate that both strategies contribute positively to the overall reasoning quality of step-level tree search methods.
@@ 4.1 Experiment setting
Evaluation Benchmark.  We tested our approach on several challenging open-source mathematical benchmarks, including MATH500 Hendrycks et al. (2021), AQuA Ling et al. (2017), OlympiadBench-TO He et al. (2024) and MATHBench Liu et al. (2024a) College-level and High-level tasks. In addition, we manually collected a selection of problems from the AMC-10 and AMC-12 competitions to serve as even more challenging benchmarks. * The AMC test set is available at https://github.com/beichenzbc/BoostStepTo simulate benchmarks with lower similarity to the example problem bank, we also conducted tests on MathVision Wang et al. (2024) and MathVerse Zhang et al. (2025), highly challenging multi-modal math benchmarks
Example Problem Bank. The example problem bank is obtained from PRM800K Lightman et al. (2023) and the steps are divided by GPT-4o.
@@
Hyper-Parameters. The temperature value is 0 in all the experiments except for step-level tree search, which needs some random sampling to generate different reasoning candidates, and the temperature value for tree search methods is set at 0.3. The reference rejection threshold is 0.7.
Prompt. Apart from some necessary guidance like step-level reasoning, we ensured that the prompts for each method were as similar as possible to make the comparison fairer. The specific prompts are listed in the supplementary materials.
@@ 4.2 Comparing to Problem-Level ICL
Table 1: A comparison of different in-context learning strategies on different benchmarks on GPT-4o and Qwen2.5-Math-72B-Instruct. The example problem bank is constructed from PRM800K, so MATH500 is an in-domain benchmark while others are all out-domain benchmarks. The best results are in bold.
@@
We conduct a rigorous comparison of our step-level in-context learning and traditional problem-level few-shot learning in various aspects. For traditional problem-level few-shot learning, we set the shot number to 4, which is a common setting.
Better Performance. We compare the performance between traditional few-shot learning and our step-level in-context learning across multiple benchmarks and base models. The results are presented in Tab. 1. Our step-level in-context learning achieves a general and significant improvement across various benchmarks compared to problem-level few-shot learning.
@@
few-shot 28.7 (-1.9) 53.2 (0.0)
Ours 35.2 (+4.6) 54.2 (+1.0)
Table 3: Experiments on the sensitivity of the similarity between the question and the example problem bank. R_t indicates that the examples are the t_th similar for different methods without any rejection strategy. Given a less similar example, our method suffers a 2.26% performance loss, which is much lower compared to few-shot learning(4.4%).
@@
Better Generalizability. Traditional few-shot learning requires the example problem bank highly similar to the questions to be solved, which limits its generalizability. To compare the generalizability, We also test different methods on multi-modal mathematical benchmarks including MathVision Wang et al. (2024) and MathVerse Zhang et al. (2025), which has much lower similarity with our example problem bank. The results are shown in Tab. 2. Problem-level few-shot learning not only fails to enhance reasoning performance but can also have a negative impact, while our method continues to achieve appreciable improvements, demonstrating better general applicability.
Better Robustness. We also manually decrease the similarity between the examples and the problems by selecting the t_th similar example during reasoning to evaluate the robustness. The result is shown in Tab. 3. We can observe that traditional problem-level in-context learning suffers from a severe decrease and is even worse than 0-shot learning when t is larger than 4. In contrast, our method does not show a significant decline and is consistently better than the 0-shot reasoning.
@@ 4.3 Construction of Example Problem Bank
To better align with the steps in reasoning, we propose constructing a step-level question bank based on the reasoning content rather than grammatical divisions. To prove our assumption, we compare our approach with a commonly used strategy that constructs steps based on grammatical segmentation, using periods ’.’ as the delimiter, on the same dataset PRM800K and under identical conditions. Results are presented in Tab. 4. Our method largely outperforms those using periods as a delimiter.
@@ 4.4 Comparison of Retrieving Strategies
The key factor in determining the effectiveness of in-context learning lies in the relevance of the retrieved examples. At the finer-grained step level, designing an appropriate retrieval strategy becomes even more crucial and challenging. Therefore, we propose the first-try strategy, which involves understanding what the model currently needs to reason about using a first attempt and then searching the problem set for similar steps to guide the model in fully outputting the current step. To validate the effectiveness of this method, we compare it with several other strategies mentioned in Sec.3.3, retrieving by the entire reasoning path $s_{i-1},s_{i-2},\ldots,s_{1},q$ or only by the immediately preceding step $s_{i-1}$.
@@
Table 5: Comparison of different retrieval strategies in step-level in-context learning. The base model is GPT-4o and all the prompts are the same. ’Path’ represents retrieving by the reasoning path including all previous steps $s_{i-1},s_{i-2},\ldots,s_{1}$ and question $q$, while ’Pre-Step’ represents retrieving by only the immediately preceding step $s_{i-1}$. The best results are in bold.
@@
4.5 Example-guided Step-level Tree Search
Table 6: A detailed ablation on incorporating retrieving similar steps to provide fine-grained guidance during the reasoning and verifying phases of step-level tree search methods. Base models are GPT-4o and prompts are the same. The best results are in bold.
@@ 4.5 Example-guided Step-level Tree Search
As discussed above, the reasoning capability of the generating model and the verifying capability of the critic model are two core factors of Step-level Tree Search methods, and our strategy can bring benefits in both ways. On one hand, it can improve the accuracy of generating candidate nodes using the previously mentioned first-try strategy when reasoning nodes are generated. On the other hand, it can increase the accuracy of evaluation by introducing similar examples during critic model assessments and therefore ensures that the correct reasoning nodes are more likely to be preserved. These two strategies can be decoupled, allowing us to demonstrate the effectiveness of each component through ablation studies.
In the setup for tree search methods, we utilize GPT-4o as the reasoning model and GPT-4o-mini as the Process-supervised Reward Model. Moreover, we adopted the Pairwise Preference Reward Model (PPRM) configuration Zhang et al. (2024b)which reduces variability and ensures a more robust evaluation. Detailed settings will be listed in the appendix.
Tab. 6 presents the results of integrating in-context learning into the reasoning and evaluation phases of Tree Search methods. The results indicate that introducing examples enhances the verifying and reasoning capabilities of tree search methods, therefore both approaches contribute to the improvement of overall reasoning performance by providing similar ground truth steps. Specifically, introducing example steps in reasoning phases will bring more benefits to the overall reasoning quality.
@@ 4.6 Case Study
Figure 4: A specific example of adjusting reasoning during real-time inference through step-level in-context learning. The first try uses a wrong equation while the retrieving example step guides the model to use the correct equation and get the correct conclusion.
Here we demonstrate a specific example of how our step-level in-context learning boosts step-level reasoning. Given the question, we first let the model have a first try on step one. Unfortunately, because the model is unfamiliar with trigonometric functions, it makes an error on the tangent sum formula, therefore leading to a wrong step. However, we can get a rough idea of what the model wants to calculate at this step according to the first try. Then, we find a similar step that correctly leverages the tangent sum formula in the step-level example problem bank. Therefore, with the guidance provided, the model correctly applied the tangent sum formula during the second reasoning attempt and arrived at the correct answer.
@@ 5 Conclusion
We propose step-level in-context learning, which provides real-time, fine-grained guidance during the reasoning process by searching for similar steps according to the first-try reasoning attempt. This approach improves the model’s reasoning capabilities and reduces the dependency on the similarity of the example problem set, thereby increasing the generalizability of in-context learning. Moreover, our method can also enhance the reasoning and evaluation capability of Step-level Tree Search methods by introducing similar steps in reasoning and verifying phases respectively, thereby improving the overall reasoning correctness.
@@ A.3 Benchmarks
We tested our approach on several mathematical benchmarks, including MATH500 Hendrycks et al. (2021), AQuA Ling et al. (2017), OlympiadBench-TO He et al. (2024) and MATHBench Liu et al. (2024a). Specifically, we use the Olympiad-TO (text-only) subset of OlympiadBench and the application problems in college-level and high-level difficulty of MATHBench.
@@
Appendix B Detailed Setup for Step-Level Tree Search
In the setup for tree search methods, we utilize GPT-4o as the reasoning model and employ GPT-4o-mini as the Process-supervised Reward Model (PRM). For the PRM, we adopted the Pairwise Preference Reward Model (PPRM) configuration Zhang et al. (2024b). Specifically, PPRM transforms the absolute rewards calculation into preference predictions between solutions to calculate rewards. This approach reduces the variability associated with scoring characteristics and thus leads to a more robust and consistent evaluation of different solutions.
The complete reasoning process in our experiment is as follows: we start with the target problem as the root node and obtain two initial solution steps through sampling to serve as the two initial parent nodes. In each step-level reasoning phase, we expand these two parent nodes through sampling, generating four candidate child nodes. Using the PPRM, we select the two child nodes with higher confidence to become the parent nodes for the next step of reasoning. This process continues until both candidate nodes have completed their reasoning paths, resulting in the final answers. Finally, PPRM is used to select the ultimate answer from these two reasoning paths.